Operations Report – November 12th

Crew 215 Operations Report 12-Nov-2019
SOL: 1
Name of person filing report: Shane Usher
Non-nominal systems: Opportunity, Suit 5 Helmet
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is repaired, Suit 5 Helmet rivetts falling apart, Suit 2 Battery non-charging, tested on multiple suits and will replace battery.
Generator: run
Hours run: 12h
From what time last night: 1730
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before the generator is run at night): 73%
Diesel Reading – 53%
Station Propane Reading – 73% today.
Ethanol Free Gasoline: N/A
Water (loft tank): 33 gallons
Water Meter: 1457225 units
Water (static tank): 450 gallons
Static to Loft Pump used – no
Water in Green Hab: 293.5 gallons, 24 L (6.5 gallons) used today
Water in Science Dome: 0 gallons
Toilet tank emptied: yes
Deimos rover used: not at the hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 110.5
Beginning charge: (Before EVA): 62%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes.
Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no
Curiosity rover used: yes
Hours: 116.1
Beginning charge: (Before EVA) : 62%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Suit 2 needs replacement battery to be changed tomorrow, Suit 5 Helmet falling apart, Suit 10 fuse replaced.
Summary of Hab operations: All systems nominal apart from; Stateroom5, Edison screw lightbulb Max 13W required; Fire alarm test button not working in Hab upper level; Fire alarm test button disintegrated in GreenHab; Numerous powerpoints not functional on ground floor of hab. Water pump is very loud.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered (24L). The heater automatically came on at 5pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Hammers used and returned, 2 fuses taken, electrical tape taken and returned
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy.
Questions, concerns and requests to Mission Support: Opportunity is now usable.

Copyright © The Mars Society. All rights reserved. | Main Site